I have so many problems with word problems and I really don't know how to set up the problem please help.
Jose invested a certain amonunt of money at 8% interest and $300 more than that amount at 11%. His total yearly interest was $69.10. How much did he invest at the 8% rate?

Answer by checkley77(12844)

.08X+.11(300+X)=69.10
.08X+33+.11X=69.10
.19X=69.10-33
.19X=36.10
X=36.10/.19
X=$190 ANS. FOR THE INVESTMENT @ 8%.
300+190=490 ANS. FOR THE INVESTMENT @ 11%.
PROOF:
.08*190+.11*490=69.10
15.2+53.9=69.10
69.1=69.1